Leslie bought 3 identical phone cases for a total cost of $20.97 including a $6.99 flat rate fee for shipping. What equation could be used to find the cost for each cell phone case, c?
A.3c + 6.99 = 20.97
B.3(c + 6.99) = 20.97
C.3/c− 6.99 = 20.97
D. c =20.97+ 6.99/3

Found 2 solutions by ewatrrr, math_tutor2020:
Answer by ewatrrr(24785)

Leslie bought '3' identical phone cases for a total cost of $20.97
including a $6.99 flat rate fee for shipping.
'c' the cost of the cell phone case
What equation 'could be used' to find the cost for each cell phone case, c?
3c + 6.99 = 20.97 ****
A. is Your Answer

Then... one could solve for 'c'
c = (20.97 - 6.99)/3

Answer by math_tutor2020(3817)


c = cost of one phone case
3c = cost of three phone cases
3c+6.99 = total cost after adding on shipping
3c+6.99 = 20.97 is the equation to set up.